# A Comprehensive Survey on Multi-modal Conversational Emotion Recognition with Deep Learning

### v1

@@ 4.1.9. Data Augmentation with Generative Adversarial Networks
Multimodal emotion recognition based on adversarial learning is an advanced direction in this field, which combines the principles of adversarial learning to improve the accuracy and robustness of emotion recognition (Ren et al., 2023), (Yuan et al., 2023). Next, we introduce the existing overall process of data augmentation based on adversarial generative networks.
1) Conditional GANs Conditional Generative Adversarial Network (cGAN) (Sun et al., 2023) is a variant of GAN that introduces conditional information to more precisely control the output of the generator. The core idea of cGAN is to pass additional condition information to the generator and discriminator during the generation process, thereby generating specific types of data based on given conditions. The main advantage of cGAN is its ability to precisely control the generation process in order to generate data that meets the conditional information. The optimization goal of cGAN is defined as follows:
@@
2) Adversarial Autoencoders Adversarial Autoencoder (AAE) (Latif et al., 2020) combines the ideas of Autoencoder and GAN. The main goal of AAE is to make this encoding space more continuous and have better data generation capabilities while learning a compressed representation of data. The training objective function of AAE usually includes two parts: one is the reconstruction error of the autoencoder, which is used to ensure the quality of the encoding, and the other is the GAN loss, which is used to make the encoding distribution more continuous and closer to the real distribution. The formula is defined as follows:
@@
$\mathcal{L}_{R}^{(AAE)}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\big{\{}||\mathbf{x}-R(E(\mathbf{x}))||^{2}\big{\}}$
@@
3) Adversarial Data Augmentation Network Adversarial Data Augmentation Network (ADAN) (Wang et al., 2022) includes the following components: autoencoder $R(E(x))$, auxiliary classifier $C(E(x))$, generator $G(z,y)$ and discriminator $D(h)$. First, ADAN aims to learn a latent representation of the input data x in order to preserve the emotional information in it. Second, it attempts to ensure that the generated latent representation is consistent with the emotional information of the input data by matching the posterior distribution $p(h|z,y)$ with $p(h|x)$. Third, ADAN simultaneously strives to minimize the reconstruction error between the input data $x$ and its reconstructed version $\hat{x}$ to ensure high-quality data reconstruction. The generator $G(z,y)$ accepts a sample $z$ drawn from an M-dimensional Gaussian distribution and a one-hot encoding of the emotion label $y$ as input, and the goal is to generate samples in the latent space such that they are indistinguishable from real samples. The discriminator $D(h)$ is optimized to distinguish whether the latent vector $h$ comes from real data or from the generator.
@@
$\mathcal{L}_{R}^{(\mathrm{ADAN})}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\{||\mathbf{x}-R(E(\mathbf{x}))||^{2}\}$
$\mathcal{L}_{E}^{(\mathrm{ADAN})}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\left\{||\mathbf{x}-R(E(\mathbf{x}))||^{2}-\sum_{k=1}^{K}y_{\mathrm{emo}}^{(k)}\log C(E(\mathbf{x}))_{k}\right\}$
@@
where $\alpha$ determines the contribution of classification error to model optimization.
@@ 4.2. Sequential Context Modeling
where $\sigma$ represents activation function.
@@
Transformers are another way of modeling sequential context (Li et al., 2020; Zhu et al., 2021; Huang et al., 2020). Transformer’s long-distance modeling capabilities are far superior to recurrent neural networks like Bi-LSTM, and Transformer can achieve parallel computing. Therefore, existing research on multi-modal emotion recognition based on sequential context modeling often regards Transformer as an important technology. The implementation details of Transformer are as follows.
@@
where $Q,K,V$ represent the query vector, key vector and value vector of multi-modal features, respectively.
@@
Finally, we use sine and cosine position encoding to obtain the position information of the emotion sequence:
@@
After Transformer was proposed, many Transformer-based multi-modal conversational emotion recognition research methods were proposed to model long-distance context dependencies (Gerczuk et al., 2021; Hazarika et al., 2021). In view of the inability of previous works to model long-distance dependencies between different modal features, Yang et al. (Yang et al., 2022) proposed a multi-modal speech emotion recognition method using context Transformer, which improves the emotional representation of the current utterance by embedding contextual information. This method can adaptively learn feature fusion between modalities.
In view of the fact that existing methods cannot dynamically identify subtle changes in emotion in multi-modal features and multi-scale features in different modal features, Liu et al. (Liu et al., 2022) proposed a multi-scale self-attention fusion emotion recognition method, which uses the self-attention mechanism to extract context-related dependencies in multi-modal features. Therefore, there is potential to use Transformers to model long-distance context dependencies. This method combines bc-LSTM and a multi-head attention mechanism to achieve fine-grained emotional information mining, and uses feature-level fusion and decision-level fusion methods to experiment with cross-modal feature fusion.
@@ 4.3. Distinguishing speaker modeling
Global GRU captures the contextual semantic information of an utterance by modeling the utterance and speaker states. Each speaker state is used to memorize a speaker-specific representation of an utterance. By distinguishing the subordination relationship between speakers and utterances, it is beneficial to model the dependency relationship between speakers and utterances, thereby enhancing the semantic representation ability of context. The formula of Global GRU is defined as follows:
@@
Speakers typically reply to conversations based on contextual information from other speakers. Therefore, speaker GRU extracts the context $c_{t}$ related to the utterance $x_{t}$. The formula is defined as follows:
@@
where $W_{\beta}$ is the learnable parameters. We first calculate the attention score of the global state in the previous $t-1$ time. The attention score give higher weight to utterances related to utterance $x_{t}$. The final context vector $c_{t}$ is obtained by the dot product of the attention score $\beta$ and the global state $g_{t}$.
@@
The emotional representation et of the utterance ut is obtained by combining the speaker’s state $q_{s(ut),t}$ and the utterance $e_{t-1}$ at time $t-1$. One underlying intuition is that context has a greater impact on utterance $u_{t}$, and $e{t-1}$ integrates emotional contextual information from other parties’ states into the emotional representation et. Therefore, we use the Emotion GRU unit to model $e_{t_{1}}$, and the formula is defined as follows:
@@
The emotion representation $e_{t}$ that combines context information and speaker status information is used for the final emotion classification.
In modeling methods based on distinguishing between speakers, Ghosal et al. (Ghosal et al., 2020) proposed COSMIC, which clarifies the relationship between the speaker and the utterance, and introduces common sense knowledge to enhance the emotional understanding of the model. COSMIC can learn a variety of different prior knowledge (e.g., event relationships and causal relationships, etc.), and can distinguish speaker information and dynamically detect the speaker’s emotional changes.
@@ 4.4.1. GNN for speaker relationship modeling
GCN extends convolution operations into graph-structured data to extract structural information. GCN performs first-order neighbor information aggregation and spectral domain estimation. The formula of GCN is defined as follows:
@@
The steps to apply GCN to the field of multi-modal emotion recognition are as follows. First, each utterance is represented as a node in the graph, and edge relationships are constructed based on the context between utterances (i.e., if there is a dialogue between utterances, an edge is constructed). We then apply GCN to the constructed dialogue graph for speaker-level information extraction. Through the above process, the model can dynamically learn the correlation between sentences. According to the definition of Equation 27, our formula for aggregating surrounding contextual utterence information is deformed as follows:
@@
GAT is a variant of GCN that aggregates surrounding neighbor node features through learnable weights with an attention mechanism. GAT captures the more important node features in the graph by calculating the degree of similarity between nodes. The formula for GAT is defined as follows:
@@
(30) $H_{i}^{+(l+1)}=ReLU\left(\sum_{r\in\mathcal{R}}\sum_{j\in\mathcal{N}_{i}^{r}}\frac{1}{|\mathcal{N}_{i}^{r}|}\left(\alpha_{ij}^{(l)}W_{\theta_{1}}^{(l)}H_{j}^{+(l)}\right.\right.+\left.\alpha_{ii}^{(l)}W_{\theta_{2}}^{(l)}H_{i}^{+(l)}\right),$
The multi-modal method based on GNN is the current mainstream research, which can consider context information and speaker relationship information simultaneously (Ghosal et al., 2019). In order to jointly learn the three tasks of sequential context information, multi-modal information interaction and multi-task representation, Zhang et al. (Zhang et al., 2023a) designed a multi-modal, multi-task interactive graph attention network (M3GAT) to simultaneously model context dependencies, multi-modal emotional interactions, and speaker dependencies. M3GAT can achieve cross-modal feature interaction, capture of sequential contextual semantic information, and correlation between tasks.
@@
In view of the fact that existing graph fusion methods will cause the model to lose important semantic information and fail to eliminate redundant information in the model, Li et al. (Li et al., 2023) proposed a graph network based on cross-modal feature complementarity, which effectively extracts the speaker’s context and interaction information through multiple hypothesis spaces of the graph. This method eliminates the heterogeneity between modalities and fuses modal information by performing different message aggregation on different nodes and edge relationships in the graph, thereby extracting contextual information and speaker relationship information.
Although existing MCER methods use GCN to model conversational relationships between speakers. In particular, the most competitive methods model the dependence of conversational relations between speakers and the importance between conversational relations by using relational graph attention networks. However, existing GCN-based multimodal conversational emotion recognition methods do not consider conversational relationships and sequential information in contextual relationships. Based on the above problems, Ishiwatari et al. (Ishiwatari et al., 2020) introduced relational position coding in RGAT to provide sequence information. The specific flow chart of RGAT is shown in Fig. 10.
@@
(31) $\left.PE_{ijr}=\left\{\begin{array}[]{rl}max(-p,min(p,j-i))&r=1,wherej\in\mathcal{N}^{1}(i)\\ max(-p,min(p,j-i))&r=2,wherej\in\mathcal{N}^{2}(i)\\ max(-f,min(f,j-i))&r=3,wherej\in\mathcal{N}^{3}(i)\\ max(-f,min(f,j-i))&r=4,wherej\in\mathcal{N}^{4}(i)\end{array}\right.\right.$
where $PE_{ijr}$ represents the relative position distance between node $i$ under relationship type $r$ and its surrounding neighbor nodes $j$. The maximum relative position distance between nodes is clipped to $p$ or $4$ , which represents the context window size. $\mathcal{N}r(i)$ represents the neighborhood of node $i$ under relationship type $r$. In order to make the position encoding information learnable, FFN is used to obtain position embeddings.

### v2

@@ 4.1.8. Data Augmentation with Generative Adversarial Networks
Multimodal emotion recognition based on adversarial learning is an advanced direction in this field, which combines the principles of adversarial learning to improve the accuracy and robustness of emotion recognition (Ren et al., 2023; Yuan et al., 2023). Next, we introduce the existing overall process of data augmentation based on adversarial generative networks.
1) Conditional GANs Conditional Generative Adversarial Network (cGAN) (Sun et al., 2023) is a variant of GAN that introduces conditional information to more precisely control the output of the generator. The core idea of cGAN is to pass additional condition information to the generator and discriminator during the generation process, thereby generating specific types of data based on given conditions. The main advantage of cGAN is its ability to precisely control the generation process in order to generate data that meets the conditional information. The optimization goal of cGAN is defined as:
@@
2) Adversarial Autoencoders Adversarial Autoencoder (AAE) (Latif et al., 2020) combines the ideas of autoencoder and GAN. The main goal of AAE is to make this encoding space more continuous and have better data generation capabilities while learning a compressed representation of data. The training objective function of AAE usually includes two parts: one is the reconstruction error of the autoencoder, which ensures the quality of the encoding, and the other is the GAN loss, which makes the encoding distribution more continuous and closer to the real distribution. The formula is defined as follows:
@@
[…]
@@
3) Adversarial Data Augmentation Network Adversarial data augmentation network (ADAN) (Wang et al., 2022) includes the following components: autoencoder $R(E(x))$, auxiliary classifier $C(E(x))$, generator $G(z,y)$ and discriminator $D(h)$. First, ADAN aims to learn a latent representation of the input data $x$ to preserve the emotional information in it. Second, it attempts to ensure that the generated latent representation is consistent with the emotional information of the input data by matching the posterior distribution $p(h|z,y)$ with $p(h|x)$. Third, ADAN simultaneously strives to minimize the reconstruction error between the input data $x$ and its reconstructed version $\hat{x}$ to ensure high-quality data reconstruction. The generator $G(z,y)$ accepts a sample $z$ drawn from an M-dimensional Gaussian distribution and a one-hot encoding of the emotion label $y$ as input, and the goal is to generate samples in the latent space such that they are indistinguishable from real samples. The discriminator $D(h)$ is optimized to distinguish whether the latent vector $h$ comes from real data or from the generator.
@@
[…]
@@
where $\alpha$ determines the contribution of classification error to model optimization. ADAN improves the authenticity of fused features and the ability to align cross-modal distributions through the confrontation between the discriminator and the generator.
Generative adversarial networks are able to generate high-quality new samples that are very close to real data through an adversarial training process. This enables the model to generate new samples with greater diversity and authenticity, thereby improving the model’s generalization ability. For tasks with scarce data, especially when there are fewer samples of a specific category, GANs can be used to enhance the dataset by generating new samples, avoiding the limitations of traditional data augmentation methods on data diversity and complexity. However, the training process of GANs is often unstable, and the adversarial process between the generator and the discriminator may cause the gradient to vanish or explode, resulting in unstable quality of the generated samples.
@@ 4.2. Sequential Context Modeling
Tri-modal Hidden Markov Model is a sequence context modeling method, which relies on the previous state and can effectively capture local dependencies. For example, Morency et al. used text, video, and audio features for the task of trimodal emotion analysis, and designed a model to extract useful information in different modal features (Morency et al., 2011). After extracting multi-modal features, the three modal features are connected and input into a Hidden Markov Chain (HMM) classifier (Morency et al., 2011) to learn the emotional state of the input signal. HMM believes that the state of the current moment is only related to the information of the previous moment, which enables the model to use the context information of the utterance. The formula of HMM is defined as follows:
$\times P(\lambda_{i}^{a},\lambda_{j}^{v},\lambda_{k}^{t}|x^{a},x^{v},x^{t})$
(17) $P(w|\lambda^{a}_{i},\lambda^{v}_{j},\lambda^{t}_{k},x^{a},x^{v},x^{t})$
$\times P(\hat{w}_{b}|\lambda_{i}^{a},\lambda_{j}^{v},\lambda_{k}^{t},x^{a},x^{v},x^{t})$
where $P(\hat{w}_{b}|\lambda_{i}^{a},\lambda_{j}^{v},\lambda_{k}^{t},x^{a},x^{v},x^{t})$ represents the probability of predicted label. HMM captures cross-modal temporal dependencies and synchronization relationships through temporal dependency structures.
HMM can effectively process sequence data, especially when there are unobservable hidden states at each time step. By modeling the hidden state, HMM can capture the time dependency in the data. Therefore, HMM has strong capabilities in sequence modeling. However, HMM is based on linear models and cannot effectively handle complex nonlinear relationships in sequence data. For nonlinear sequence data, the performance of HMM may be limited. In addition, when processing large-scale data, the computational complexity of HMM training is high, especially when the parameter space is large, computational bottlenecks may be encountered.
@@
where $\sigma$ represents activation function. LSTM extracts dynamic information within the modality by capturing time dependencies.
LSTM can remember information over long time spans and is particularly suitable for modeling long-term dependencies. However, LSTM still performs worse than other more advanced models (such as Transformer) in some extremely long sequence tasks. LSTM’s ability to retain memory over long time spans is also limited. Therefore, LSTM is suitable for most tasks that need to capture time dependencies, but for very long dependencies or large-scale data, modern models such as Transformer may provide better performance.
@@
Transformers are another way of modeling sequential context (Li et al., 2020; Zhu et al., 2021; Huang et al., 2020). Transformer’s long-distance modeling capabilities are far superior to recurrent neural networks, and Transformer can achieve parallel computing. Therefore, existing research on multi-modal emotion recognition based on sequential context modeling often regards Transformer as an important technology. The implementation details of Transformer are as follows.
@@
where $Q,K,V$ represent the query vector, key vector and value vector of multi-modal features, respectively. Transformer automatically captures key modal and key area information by dynamically assigning weights.
@@
Finally, position encoding is used to obtain the position information of the emotion sequence:
@@
Transformer uses a self-attention mechanism to process input data, which is independent of the order of the sequence. Compared with traditional RNNs and LSTMs, Transformer can calculate the relationship between each input position in parallel, which significantly speeds up the training process. Due to its parallelism, Transformer is more suitable for training large-scale datasets and can greatly reduce training time. In addition, Transformer’s Multi-head Self-Attention allows the model to focus on different parts of the input from multiple perspectives, which helps capture multi-level information. However, since Transformer involves interactions with all positions when calculating self-attention, its time complexity and space complexity are both high, especially when processing long sequences, the amount of calculation increases quadratically. For long sequences, Transformer’s computational and memory requirements are very large. In addition, when the amount of data is limited, Transformer may not be able to fully exert its advantages, requiring a large amount of pre-training data and careful tuning.
After Transformer was proposed, many Transformer-based multi-modal conversational emotion recognition methods were proposed to model long-distance context dependencies (Gerczuk et al., 2021; Hazarika et al., 2021). Previous works failed to model long-distance dependencies between different modal features. To address this, Yang et al. (Yang et al., 2022) proposed a multimodal speech emotion recognition method using a context Transformer, which enhances the emotional representation of the current utterance by embedding contextual information. This method can adaptively learn feature fusion between modalities.
Existing methods struggle to dynamically identify subtle emotional changes in multimodal and multi-scale features. To address this, Liu et al. (Liu et al., 2022) proposed a multi-scale self-attention fusion emotion recognition method, which uses the self-attention mechanism to extract context-related dependencies in multimodal features. Therefore, there is potential to use Transformers to model long-distance context dependencies. This method combines bc-LSTM and a multi-head attention mechanism to achieve fine-grained emotional information mining, and uses feature-level fusion and decision-level fusion methods to experiment with cross-modal feature fusion.
@@ 4.3. Distinguishing speaker modeling
Global GRU captures the contextual semantic information of an utterance by modeling the utterance and speaker states. Each speaker state is used to memorize a speaker-specific representation of an utterance. By distinguishing the subordination relationship between speakers and utterances, it is beneficial to model the dependency relationship between speakers and utterances, thereby enhancing the semantic representation ability of context. The formula of Global GRU is defined as:
@@
Speakers typically reply to conversations based on contextual information from other. Therefore, speaker GRU extracts the context $c_{t}$ related to the utterance $x_{t}$. The formula is defined as follows:
@@
where $W_{\beta}$ is the learnable parameters. First, calculate the attention score of the global state in the previous t1t-1 time. The attention score assigns higher weight to utterances related to utterance $x_{t}$. The final context vector $c_{t}$ is obtained by the dot product of the attention score $\beta$ and the global state $g_{t}$.
@@
The emotional representation et of the utterance $u_{t}$ is obtained by combining the speaker’s state $q_{s(ut),t}$ and the utterance $e_{t-1}$ at time t1t-1. One underlying intuition is that context has a greater impact on utterance $u_{t}$, and $e_{t-1}$ integrates emotional contextual information from other parties’ states into the emotional representation $e_{t}$. Therefore, we use the Emotion GRU unit to model $e_{t_{1}}$, and the formula is defined as follows:
@@
The emotion representation $e_{t}$ that combines context information and speaker status information is used for the final emotion classification. Distinguishing speaker modeling method combines global context with individual information to improve understanding of complex information such as semantics and emotions.
Similar to traditional RNN and LSTM models, distinguishing speakers modeling methods have the ability to capture long-range dependencies. In multi-turn conversations, the previous conversation content often affects subsequent understanding and generation. Distinguishing speakers modeling methods can effectively learn these long-term dependencies and help understand contextual information. However, in scenarios where the amount of data is small or the conversation data is simple, distinguishing speakers modeling methods may suffer from overfitting. In particular, when there are not many conversation rounds in the training data set, the model tends to remember specific patterns in the training set and cannot effectively generalize to new data.
In modeling methods based on distinguishing between speakers, Ghosal et al. (Ghosal et al., 2020) proposed commonsense knowledge for emotion identification in conversations (COSMIC), which clarifies the relationship between the speaker and the utterance. It also introduces common sense knowledge to enhance the emotional understanding of the model. COSMIC can learn a variety of different prior knowledge (e.g., event relationships and causal relationships, etc.), and can distinguish speaker information and dynamically detect the speaker’s emotional changes.
@@ 4.4.1. GNN for speaker relationship modeling
Graph convolutional network (GCN) extends convolution operations into graph-structured data to extract structural information. GCN performs first-order neighbor information aggregation and spectral domain estimation. The formula of GCN is defined as follows:
@@
The steps to apply GCN to the field of multi-modal emotion recognition are as follows. First, each utterance is represented as a node in the graph, and edge relationships are constructed based on the context between utterances. We then apply GCN to the constructed dialogue graph for speaker-level information extraction. Through the above process, the model can dynamically learn the correlation between sentences. According to the definition of Equation 27, our formula for aggregating surrounding contextual utterence information is deformed as follows:
@@
The multi-layer convolutional structure of GCN can effectively integrate feature information from different modalities and enhance the expressiveness of emotional information. Especially when the amount of information is large, GCN can effectively aggregate the features of each modality, which helps the accuracy of emotional classification or regression tasks. However, GCN has difficulties in modeling long-distance dependencies. As the depth of the graph increases, the problem of over-smoothing will occur during the information propagation process, causing the representation of nodes to become similar and lose their original distinguishability.
Graph Attention Network (GAT) is a variant of GCN that aggregates surrounding neighbor node features through learnable weights with an attention mechanism. GAT captures the more important node features in the graph by calculating the degree of similarity between nodes. The formula for GAT is defined as follows:
@@
(30) $H_{i}^{+(l+1)}=ReLU\left(\sum_{r\in\mathcal{R}}\sum_{j\in\mathcal{N}_{i}^{r}}\frac{1}{|\mathcal{N}_{i}^{r}|}\left(\alpha_{ij}^{(l)}W_{\theta_{1}}^{(l)}H_{j}^{+(l)}\right.\right.+\left.\alpha_{ii}^{(l)}W_{\theta_{2}}^{(l)}H_{i}^{+(l)}\right)$
GAT dynamically distinguishes the importance of neighboring nodes, giving the model stronger nonlinear expression and local pattern capture capabilities.
In multimodal emotion recognition, the relationship between modalities may be heterogeneous, that is, the data of different modalities may be different in nature. GAT can effectively handle such heterogeneous graphs because it gives different weights to each type of node through the attention mechanism, thereby better representing the heterogeneous relationship between different modalities. However, although the attention mechanism can improve the flexibility of information aggregation, it may also lead to excessive focus on local information, especially in some tasks that need to consider the global context. The local weighting of GAT may limit the global learning ability of the model.
The multi-modal method based on GNN is the current mainstream research, which can consider context information and speaker relationship information simultaneously (Ghosal et al., 2019). To jointly learn sequential context, multimodal interaction, and multitask representation, Zhang et al. (Zhang et al., 2023a) designed the M3GAT (Multi-modal, Multi-task Interactive Graph Attention Network). M3GAT simultaneously models context dependencies, multimodal emotional interactions, and speaker dependencies. It enables cross-modal feature interaction, captures sequential contextual semantic information, and establishes task correlations.
@@
Existing graph fusion methods often cause the model to lose important semantic information and fail to eliminate redundancy. To address this, Li et al. (Li et al., 2023) proposed a graph network based on cross-modal feature complementarity. This method effectively extracts the speaker’s context and interaction information using multiple hypothesis spaces in the graph. This method eliminates the heterogeneity between modalities and fuses modal information by performing different message aggregation on different nodes and edge relationships in the graph, thereby extracting contextual information and speaker relationship information.
Although existing MCER methods use GCN to model conversational relationships between speakers. In particular, the most competitive methods model the dependence of conversational relations between speakers and the importance between conversational relations by using relational graph attention networks. However, existing GCN-based multimodal conversational emotion recognition methods do not consider conversational relationships and sequential information in contextual relationships. Based on the above problems, Ishiwatari et al. (Ishiwatari et al., 2020) introduced relational position coding in relationship graph attention network (RGAT) to provide sequence information. The specific flow chart of RGAT is shown in Fig. 12.
@@
[…]
where $PE_{ijr}$ represents the relative position distance between node $i$ under relationship type $r$ and its surrounding neighbor nodes $j$. The maximum relative position distance between nodes is clipped to $p$ or $4$ , which represents the context window size. $\mathcal{N}r(i)$ represents the neighborhood of node $i$ under relationship type $r$. To make the position encoding information learnable, feedforward network (FFN) is used to obtain position embeddings.
RGAT improves emotion recognition performance by explicitly modeling complex relationships within and between modalities. However, if the prior relationship definition is unreasonable (e.g., incorrectly connecting irrelevant modalities), the performance may deteriorate.